use the five step problem solving process
translate
reword
carry-out
check
state
for a holiday party marcus is making a five layer torte which consist of 5-9 inch diameter cake layers with a cream topping on each layer. he knows that one recipe of cream will cover 60 square inches of cake. how many batches of the topping must he make to complete his torte? (use n=3.14 as an approximate value)

Answer by stanbon(75887)

for a holiday party marcus is making a five layer torte which consist of 5-9 inch diameter cake layers with a cream topping on each layer. he knows that one recipe of cream will cover 60 square inches of cake. how many batches of the topping must he make to complete his torte? (use n=3.14 as an approximate value)
-----------------
area of one of the layers: (pi)4.5^2 = 63.59 sq. in.
---
area of 5 of these layers: = 5*63.59 = 317.93 sq. in.
----
# of batches = 317.93/60 = 5.3
-----
Rounding up you get # of batches is 6.
=========================================
Cheers,
Stan H.